# Temporal Triplane Transformers as Occupancy World Models

### v1

@@ Abstract
Recent years have seen significant advances in world models, which primarily focus on learning fine-grained correlations between an agent’s motion trajectory and the resulting changes in its surrounding environment. However, existing methods often struggle to capture such fine-grained correlations and achieve real-time predictions. To address this, we propose a new 4D occupancy world model for autonomous driving, termed T3Former. T3Former begins by pre-training a compact triplane representation that efficiently compresses the 3D semantically occupied environment. Next, T3Former extracts multi-scale temporal motion features from the historical triplane and employs an autoregressive approach to iteratively predict the next triplane changes. Finally, T3Former combines the triplane changes with the previous ones to decode them into future occupancy results and ego-motion trajectories. Experimental results demonstrate the superiority of T3Former, achieving 1.44$\times$ faster inference speed (26 FPS), while improving the mean IoU to 36.09 and reducing the mean absolute planning error to 1.0 meters.
Keywords Occupancy world models  $\cdot$ Triplane  $\cdot$ Multi-scale  $\cdot$ Autoregression
@@ 1 Introduction
footnotetext: Corresponding authors
World models [1, 2] are designed to predict future scenes and facilitate motion planning for agents. These models first construct lower-dimensional representations of the scenes, which serve as a foundation for learning the patterns of environmental dynamics. This capability supports the identification of potential dangers, the determination of traffic participants’ intentions, and ultimately leads to improved decision-making.
This paper focuses on world models for autonomous driving [3, 4, 5, 6, 7], where accurately predicting the future behavior of traffic participants is essential for the agent’s planning. Existing methods [8, 6, 7, 9] mainly provide instance-level predictions for traffic participants from a Bird’s Eye View (BEV) perspective, or directly utilize diffusion models [10, 11, 12, 13, 14] to generate future pixel-level driving views. However, these methods have difficulty in establishing fine-grained, 3D associations between changes in the scene and the agent’s motion planning. Recent advancements in 3D occupancy technologies [15, 16, 17, 18, 19] have gained significant attention from both academia and industry [20, 21]. 3D occupancy offers a well structured 3D representations of surroundings, which facilitates motion planning by establishing fine-grained correlations between surrounding occupied voxels and the agent’s actions.
@@
In this context, the first task is to learn a compact, lower-dimensional representation of the raw occupancy data [4, 3, 5], which must preserve both geometric and semantic details. Current methods mainly employ VQ-VAE [22] for the purpose. While VQ-VAE has shown success in several generative tasks [23, 24, 25], its application to compressing occupancy data forces the latent space of 3D structures to be discretized. This results in different 3D structures sharing the same token in the codebook, inevitably causing a loss of 3D structural features and reduced reconstruction accuracy. Furthermore, in real-world scenarios, most voxels in the occupancy data are empty (unoccupied) [26], and some objects contain only a small number of semantic voxels (e.g., pedestrians, road signs). This leads to the omission of sparse but critical details in the compressed features.
(a) Parallel forecasting.
(b) Autoregressive forecasting.
(c) Performance comparison.
Figure 1: Different principles of occupancy world models: (a) Parallel forecasting [4, 5], which inputs all future queries and historical occupancy results, using a Transformer [27] architecture to predict future occupancy results in parallel; and (b) Autoregressive forecasting [3], which predicts future results iteratively, conditioning each step on a single future query and the previous frame’s prediction. Existing models typically use pre-trained VQ-VAE [22] to compress occupancy results into a compact representative code for predicting future codes. In contrast, our T3Former, shown in (b.2), employs more precise and efficient triplane representations and leverages multi-scale Transformers to predict triplane changes. Therefore, we achieve faster and more accurate occupancy forecasting, as shown in (c).
@@
Based on the scene representation, recent world models [5, 4] often employ Transformers for future state prediction. A key challenge for these approaches is capturing the diverse motion patterns of multi-scale objects on the road. For example, small objects like pedestrians or bicycles may exhibit sudden and abrupt motion changes, while larger objects such as trucks or buses tend to demonstrate more inertial movement, potentially requiring a different modeling granularity. Furthermore, they employ the Transformer to directly predict the complete future occupancy state. This end-to-end approach requires the model to learn both state changes and absolute state reconstruction simultaneously, increasing the learning burden. This not only demands larger model capacity but also increases the risk of error accumulation in long-term predictions.
@@
We present a novel occupancy world model using temporal triplane transformers (T3Former) that addresses the aforementioned challenges. First, we propose to compress the 3D occupancy data into a triplane structure, a concept from computer graphics [28, 29, 30, 31, 32, 33]. A triplane comprises three planes – $xy$, $xz$, and $yz$ – onto which the occupancy data are projected. Compared to VQ-VAE, this representation ensures high reconstruction accuracy of the occupancy grid, achieving a 20% improvement in mIoU while reducing the latent space size by 34%. Next, we extend the representation to the temporal dimension ($xyt$, $xzt$, and $yzt$), using Transformer models across multiple scales to regressively generate incremental changes in the triplane. Finally, we leverage these changes, along with the previous occupancy state, to predict future occupancy and perform motion planning.
@@
Compared to OccWorld, T3Former reduces the total model size by 23%, while also mitigating the cumulative error in long-term occupancy predictions, resulting in a 110.6% increase in mIoU. At the same time, the motion planning accuracy achieves the lowest average error of 1.0 meters and the lowest average collision rate of 30%. This is mainly because we focus on predicting multi-scale occupancy state changes within the triplane space, which more effectively capture environmental dynamics. In summary, our contributions are mainly three-fold:
@@
  • We design a new 4D autoregressive occupancy world model, T3Former, that enhances long-term scene forecasting and enables precise motion planning.
@@
  • T3Former pre-trains a compact occupancy triplane representation to predict future incremental changes of triplanes using multi-scale Transformers. These changes, combined with the previous frame’s output, are then decoded into occupancy states and motion trajectories.
@@
  • Extensive experiments validate our state-of-the-art (SOTA) performance in terms of occupancy forecasting, motion planning, and real-time execution.
@@ 3.1 Formulation
Note that our OWM, $\bm{\Phi}=\{\bm{\Phi}_{\rm enc},\bm{\Phi}_{\rm dec},\bm{\Phi}_{\rm fut},\bm{\Phi}_{\rm act}\}$, operates in an autoregressive fashion. In particular, Equations (3) and (4) iteratively use previously predicted outcomes as part of the historical data to forecast future scenes and provide motion proposals.
(a) Pretrained traiplane for OWM.
(b) Temporal Triplane Transformers (T3Former) as OWM.
Figure 2: Workflow of T3Former. T3Former first pre-trains compact triplane representations of occupancy data. Next, T3Former utilizes multi-scale Transformers to capture comprehensive temporal dynamics within each plane, thereby predicting future changes in the triplane. Finally, T3Former leverages these predicted changes, along with the previous triplane, to generate future motion proposals.
@@ 3.2.1 Pre-training triplane representations for OWM
Specifically, as shown in Figure 2(a), the occupancy data $S^{t}$ is first encoded by $\bm{\Phi}_{\rm enc}$ to produce $s^{t}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}\times L_{s}}$. Then, an axis-wise average pooling operation is applied to obtain three orthogonal feature planes, $s^{t}=[s^{t}_{xy},s^{t}_{xz},s^{t}_{yz}]$, where $s^{t}_{xy}\in\mathbb{R}^{C_{s}\times W_{s}\times L_{s}}$, $s^{t}_{xz}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}}$, and $s^{t}_{yz}\in\mathbb{R}^{C_{s}\times H_{s}\times L_{s}}$, corresponding to the $xy$, $xz$, and $yz$ planes, respectively. To decode the original occupancy data, for each point in the 3D occupancy grid, denoted as $pos=(x,y,z)$, it serves as a position query to retrieve the corresponding features from the three planes in $s^{t}$. These features are then summed and concatenated with the positional encoding ${\rm PE}(pos)$, which is passed into $\bm{\Phi}_{\rm dec}$ to predict the semantic label for the position $pos$. The detailed structures of $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ are given in the supplementary martial.
By pretraining $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ with Equation (5), we obtain a highly generalizable and accurate triplane representation:
@@ 3.2.2 Multi-scale autoregressions in T3Former
Figure 3: Motion planning with T3Former.
In particular, $\bm{\Phi}_{\rm fut}=\{\bm{\Phi}_{{\rm fut}_{xy}},\bm{\Phi}_{{\rm fut}_{xz}},\bm{\Phi}_{{\rm fut}_{yz}}\}$. Each plane’s prediction model, $\bm{\Phi}_{{\rm fut}_{i}}$, consists of Transformers [27] operating at multiple scales. These Transformers share the same architecture but have distinct learnable parameters and varying input sizes, depending on the plane and scale, as shown in Figure 2(b). For example, in the case of $\bm{\Phi}_{{\rm fut}_{xy}}$, the input $s^{t}_{xy}$ is passed through UNet [46]-style downsampling, resulting in $V$ scales of plane features. The feature at each scale $v\in V$ is denoted as $s^{t,v}_{xy}\in\mathbb{R}^{C_{s_{xy}}^{v}\times W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}}$. These features are then flattened across the last two dimensions to form $W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}$ tokens, which are input into the Transformer encoder to generate spatio-temporal memory. Next, we employ a learnable query $Q^{t,v}_{s_{xy}}$, which has the same dimension as $s^{t,v}_{xy}$, to enable cross-attention in the Transformer decoder and generate the token change from the current to the next time step based on the memory. Finally, the token changes at all scales are aggregated using UNet-style upsampling, yielding the plane’s feature change at the next time step, $\Delta\hat{s}^{t+1}_{xy}$. This change is then combined with the previous plane state, $\hat{s}^{t}_{xy}$, through a convolutional operation $\phi$ with a $1\times 1$ kernel size.
@@
Since we have pre-trained the triplane occupancy representations, the future triplane can be used as the ground truth (GT) to supervise T3Former. Besides, we finetune $\Phi_{\rm dec}$ to the separate predictions of each plane. Formally:
@@ 3.3 Motion planning with T3Former
4 Experiments
4.1 Experimental settings
Targets and evaluation metrics. The main focus of OWM is jointly modeling occupancy forecasting and motion planning. Following the conventions in [3, 5, 4], we use the past four frames (2 seconds) to predict the outcomes of the next six frames (3 seconds). We conduct two sets of experiments:
  • To validate the accuracy of 4D occupancy forecasting, we use intersection over union (IoU) and mean IoU (mIoU) across all semantic classes to measure the predicted future occupancy under the Occ3D dataset [17].
  • To assess the precision and safety of motion planning, we calculate the L2 distance (in meters) between the planned and GT trajectories, and the collision rate with traffic participants’ bounding boxes using the nuScenes dataset [49].
@@ 4.1 Experimental settings
Implementation details. The dataset consists of 1,000 scenes, of which 850 are used for training and 100 for testing. Each scene contains up to 40 timesteps, with a sampling frequency of 2Hz. The dimensions of the occupancy data $S^{t}$ at each timestep are $16\times 200\times 200$. The pre-trained Triplane feature dimensions are $8\times 16\times 100\times 100$. In $\bm{\Phi}_{\rm fut}$, predictions for each plane incorporate features at $V=5$ different scales, while the token dimension in $\bm{\Phi}_{\rm act}$ is $d_{\rm act}=50$. The objectives, $J_{\rm enc,dec}$, $J_{\rm fut}$, and $J_{\rm mot}$, are optimized using AdamW with a weight regularization factor of 0.01, an initial learning rate of 0.001, and cosine decay with a minimum learning rate of $10^{-6}$. We first pre-train $\bm{\Phi}_{\rm enc}$ and $\Phi_{\rm dec}$ with a batch size of 10, using random flip augmentation to obtain the triplane representations. Then, we train $\Phi_{\rm fut}$ and $\bm{\Phi}_{\rm act}$ with a batch size of 1, while fine-tuning $\bm{\Phi}_{\rm dec}$. All training and testing are performed on 4 RTX 4090 GPUs. More details are provided in the supplementary material.
4.2 Comparisons with the state-of-the-art
4.2.1 4D occupancy forecasting
Table 1 presents the performance results of various methods, considering two modes of operation: (i) using 3D occupancy GTs as historical input, marked with “-O”; (ii) using predicted 3D occupancy data from FB-Occ [48] as historical input, marked with “-F”. “Copy&Paste” refers to directly using the GT occupancy data from the current timestep as predictions for future outcomes. Vanilla-LLaMA-O refers to the method that simply flattens the occupancy data and feeds it into LLaMA [54] for training.
4.2.2 Motion planning
We extensively compare T3Former with SOTA methods for autonomous driving, including LiDAR-based methods (IL [50], NMP [51], FF [52], and EO [53]), camera-based methods (ST-P3 [8], UniAD [6], VAD [7], and OccNet [15]), and occupancy-based methods (OccNet [15], OccWorld [3], and OccLLaMA [4]). LiDAR- and camera-based methods explore various combinations of auxiliary supervisions to enhance the quality of planned motion trajectories. Despite their success, the auxiliary information they rely on is labor-intensive to obtain.
In contrast, occupancy-based methods generate motion proposals using only occupancy representations. T3Former achieves the lowest L2 error over time, the lowest collision rate at the final time step, and the lowest average collision rate across all future motions. Compared to SOTA occupancy-based methods, T3Former reduces the average L2 error by 12.3% and the average collision rate by 38.8%. This is primarily because T3Former can establish precise, task-relevant correlations between scene changes and motion trajectories in a deep latent space, effectively filtering out noise from scene variations.
Figure 4: Visualization of 4D occupancy forecasting and motion planning for the next 3 seconds (zoom in for a clearer view).

### v2

@@ Abstract
World models aim to learn or construct representations of the environment that enable the prediction of future scenes, thereby supporting intelligent motion planning. However, existing models often struggle to produce fine-grained predictions and to operate in real time. In this work, we propose T3Former, a novel 4D occupancy world model for autonomous driving. T3Former begins by pre-training a compact triplane representation that efficiently encodes 3D occupancy. It then extracts multi-scale temporal motion features from historical triplanes and employs an autoregressive approach to iteratively predict future triplane changes. Finally, these triplane changes are combined with previous states to decode future occupancy and ego-motion trajectories. Experimental results show that T3Former achieves 1.44$\times$ speedup (26 FPS), improves mean IoU to 36.09, and reduces mean absolute planning error to 1.0 meters. Demos are available in the supplementary material.
@@ 1 Introduction
World models [10, 9] aim to represent the environment, predict future scenes and enable agents to perform advanced motion planning. These models first encode sensory inputs into compact, lower-dimensional representations, which are then used to model the temporal evolution of the environment. When applied to autonomous driving [58, 46, 51, 16, 8, 53], world models help anticipate potential hazards and support more informed, safer decision-making [6].
Existing models [14, 16, 19, 57] focus on predicting future instance segmentation of traffic participants from a Bird’s Eye View (BEV), or use diffusion models [48, 45, 44, 7, 18] to generate future driving scenes. However, these approaches encounter challenges in establishing fine-grained, 3D associations between scene changes and the agent’s motion planning. Recently, 3D occupancy  [40, 32, 39, 47, 28] has emerged as a structured and spatially rich representation of the environment. It enables finer correlations between occupied voxels and planning decisions, thus offering improved support for downstream control tasks. In this paper, we focus on the design of Occupancy World Models (OWMs) to address this challenge.
(a) Conventional approach.
(b) Our approach T3Former.
Figure 1: Different occupancy world architectures: Conventional approaches [58, 46, 51, 53, 8], as shown in (a), use a pre-trained VAE [41, 24] to compress occupancy data (denoted by $S$) into a compact representation code, which is then combined with encoded previous actions (marked as $a$) to predict future codes and actions through the Transformer. In contrast, our method, T3Former, shown in (b), employs more precise and efficient triplane representations and leverages a lightweight Transformer model to separately predict the triplane changes of each plane ($\Delta$). These $\Delta$ values are then used to for occupancy forecasting and motion planning.
@@
In this context, the first task is to learn a compact representation of the raw occupancy data [46, 58, 51], which must preserve both geometric and semantic details. Current methods mainly employ VQ-VAE [41, 24] for the purpose. While VQ-VAE has shown success in several generative tasks [35, 20, 55], its application to compressing occupancy data forces the latent space of 3D structures to be discretized. This results in different 3D structures sharing the same token in the codebook, inevitably causing a loss of 3D structural features and reduced accuracy. Furthermore, in real-world scenarios, most voxels in the occupancy data are empty (unoccupied) [43], and some objects contain only a small number of semantic voxels (e.g., pedestrians, road signs). This leads to the omission of sparse but critical details in the compressed features.
Figure 2: Performance comparisons on the test dataset. Our T3Former achieves more accurate occupancy forecasting and motion planning with faster inference speed.
@@
Based on the scene representation, recent world models [51, 46] often use Transformers for future state prediction. A key challenge here is capturing the diverse motion patterns of multi-scale objects on the road. For example, small objects like pedestrians or bicycles may exhibit sudden and abrupt motion changes, while larger objects such as trucks or buses tend to demonstrate more inertial movement, potentially requiring a different modeling granularity. Furthermore, they employ the Transformer to directly predict the complete future occupancy state. This end-to-end approach requires the model to learn both state changes and absolute state reconstruction simultaneously, increasing the learning burden. This not only demands larger model capacity but also increases the risk of error accumulation in long-term predictions.
@@
We present a novel occupancy world model using temporal triplane transformers, termed T3Former, that addresses the aforementioned challenges. First, we propose to compress the 3D occupancy data into a triplane structure, a concept from computer graphics [21, 37, 15, 38, 17, 25]. A triplane comprises three planes – $xy$, $xz$, and $yz$ – onto which the occupancy data are projected. Compared to VQ-VAE, this representation ensures high reconstruction accuracy of the occupancy grid, achieving a 20% improvement in mIoU while reducing the latent space size by 34%. Next, we extend the representation to the temporal dimension ($xyt$, $xzt$, and $yzt$), using Transformer models across multiple scales to regressively generate incremental changes in the triplane. Finally, we leverage these changes, along with the previous occupancy state, to predict future occupancy and perform motion planning.
@@
Compared to conventional approaches, we focus on predicting occupancy state changes across varying scales within the triplane space, which more effectively capture environmental dynamics. As a result, as shown in Figure 2, T3Former mitigates cumulative errors in long-term occupancy predictions, leading to an 81.1% increase in averaged mIoU. Additionally, motion planning accuracy achieves the lowest average error of 1.0 meter and the lowest average collision rate of 30%. Furthermore, by leveraging the concept of $\Delta$ predictions, T3Former significantly reduces model capacity, achieving 26 FPS. In summary, our contributions are mainly three-fold:
@@
$\bullet$ We design a novel 4D autoregressive occupancy world model, T3Former, that enhances long-term scene forecasting and enables precise motion planning.
@@
$\bullet$ T3Former pre-trains a compact occupancy triplane representation to predict future incremental changes of triplanes using multi-scale Transformers. These changes, combined with the previous frame’s output, are then decoded into occupancy states and motion trajectories.
@@
$\bullet$ Extensive experiments validate our state-of-the-art (SOTA) performance in terms of occupancy forecasting, motion planning, and real-time execution.
@@ 3.1 Formulation
Note that our OWM, $\bm{\Phi}=\{\bm{\Phi}_{\rm enc},\bm{\Phi}_{\rm dec},\bm{\Phi}_{\rm fut},\bm{\Phi}_{\rm act}\}$, operates in an autoregressive fashion. In particular, Eqs. (3) and (4) iteratively use previously predicted outcomes as part of the historical data to forecast future scenes and provide motion proposals. Figure 3 illustrates our design, detailed below.
Figure 3: Workflow of T3Former. T3Former first pre-trains compact triplane representations of occupancy data. Next, T3Former utilizes multi-scale Transformers to capture comprehensive temporal dynamics within each plane, thereby predicting future triplane changes. Finally, T3Former leverages these predicted changes, along with the previous triplane, to generate future motion proposals.
@@ 3.2.1 Pre-training triplane representations for OWM
Specifically, as shown in Figure 4, the occupancy data $S^{t}$ is first encoded by $\bm{\Phi}_{\rm enc}$ to produce $s^{t}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}\times L_{s}}$. Then, an axis-wise average pooling operation is applied to obtain three orthogonal feature planes, $s^{t}=[s^{t}_{xy},s^{t}_{xz},s^{t}_{yz}]$, where $s^{t}_{xy}\in\mathbb{R}^{C_{s}\times W_{s}\times L_{s}}$, $s^{t}_{xz}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}}$, and $s^{t}_{yz}\in\mathbb{R}^{C_{s}\times H_{s}\times L_{s}}$, corresponding to the $xy$, $xz$, and $yz$ planes, respectively. To decode the original occupancy data, for each point in the 3D occupancy grid, denoted as $pos=(x,y,z)$, it serves as a position query to retrieve the corresponding features from the three planes in $s^{t}$. These features are then summed and concatenated with the positional encoding ${\rm PE}(pos)$, which is passed into $\bm{\Phi}_{\rm dec}$ to predict the semantic label for the position $pos$. The detailed structures of $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ are given in the supplementary martial.
By pretraining $\bm{\Phi}_{\rm enc}$ and $\bm{\Phi}_{\rm dec}$ with Eq. (5), we obtain a highly generalizable and accurate triplane representation:
@@ 3.2.2 Multi-scale autoregressions in T3Former
In particular, $\bm{\Phi}_{\rm fut}=\{\bm{\Phi}_{{\rm fut}_{xy}},\bm{\Phi}_{{\rm fut}_{xz}},\bm{\Phi}_{{\rm fut}_{yz}}\}$. Each plane’s prediction model, $\bm{\Phi}_{{\rm fut}_{i}}$, consists of Transformers [42] operating at multiple scales. These Transformers share the same architecture but have distinct learnable parameters and varying input sizes, depending on the plane and scale, as shown in Figure 3. For example, in the case of $\bm{\Phi}_{{\rm fut}_{xy}}$, the input $s^{t}_{xy}$ is passed through UNet [36]-style downsampling, resulting in $V$ scales of plane features. The feature at each scale $v\in V$ is denoted as $s^{t,v}_{xy}\in\mathbb{R}^{C_{s_{xy}}^{v}\times W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}}$. These features are then flattened across the last two dimensions to form $W_{s_{xy}}^{v}\times H_{s_{xy}}^{v}$ tokens, which are input into the Transformer encoder to generate spatio-temporal memory. Next, we employ a learnable query $Q^{t,v}_{s_{xy}}$, which has the same dimension as $s^{t,v}_{xy}$, to enable cross-attention in the Transformer decoder and generate the token change from the current to the next time step based on the memory. Finally, the token changes at all scales are aggregated using UNet-style upsampling, yielding the plane’s feature change at the next time step, $\Delta\hat{s}^{t+1}_{xy}$. This change is then combined with the previous plane state, $\hat{s}^{t}_{xy}$, through a convolutional operation $\phi$ with a $1\times 1$ kernel size.
@@
Since we have pre-trained the triplane occupancy representations, the future triplane can be used as the ground truth (GT) to supervise T3Former. Additionally, as we predict each plane independently, misalignments may occur. We finetune $\Phi_{\rm dec}$ to the align predictions of each plane. Formally:
@@ 3.3 Motion planning with T3Former
4 Experiments
4.1 Experimental settings
Targets and evaluation metrics. OWMs focus on jointly modeling occupancy forecasting and motion planning. Following the conventions in [58, 51, 46], we use the past four frames (2 seconds) to predict the outcomes of the next six frames (3 seconds). We conduct two sets of experiments:
$\bullet$ To validate the accuracy of 4D occupancy forecasting, we use intersection over union (IoU) and mean IoU (mIoU) across all semantic classes to measure the predicted future occupancy under the Occ3D dataset [39].
$\bullet$ To assess the precision and safety of motion planning, we calculate the L2 distance (in meters) between the planned and GT trajectories, and the collision rate with traffic participants’ bounding boxes using the nuScenes dataset [3].
Implementation details. The dataset consists of 1,000 scenes, of which 850 are used for training and 100 for testing. Each scene contains up to 40 timesteps, with a sampling frequency of 2Hz. The dimensions of the occupancy data $S^{t}$ at each timestep are $16\times 200\times 200$. The pre-trained Triplane shape is $16\times 100\times 100$ with 8 channels. In $\bm{\Phi}_{\rm fut}$, predictions for each plane incorporate features at $V=5$ different scales, while the token dimension in $\bm{\Phi}_{\rm act}$ is $d_{\rm act}=50$. The objectives, $J_{\rm enc,dec}$, $J_{\rm fut}$, and $J_{\rm mot}$, are optimized using AdamW with a weight regularization factor of 0.01, an initial learning rate of 0.001, and cosine decay with a minimum learning rate of $10^{-6}$. We first pre-train $\bm{\Phi}_{\rm enc}$ and $\Phi_{\rm dec}$ with a batch size of 10, using random flip augmentation to obtain the triplane representations. Then, we train $\Phi_{\rm fut}$ and $\bm{\Phi}_{\rm act}$ with a batch size of 1, while fine-tuning $\bm{\Phi}_{\rm dec}$. All training and testing are performed on 4 RTX 4090 GPUs. More details are provided in the supplementary material.
@@ 4.1 Experimental settings
4.2 Comparisons with the state-of-the-art
4.2.1 4D occupancy forecasting
Table 1 presents the performance results of various methods, considering two modes of operation: (i) using 3D occupancy GTs as historical input, marked with “-O”; (ii) using predicted 3D occupancy data from FB-Occ [27] as historical input, marked with “-F”. “Copy&Paste” refers to directly using the GT occupancy data from the current timestep as predictions for future outcomes.
4.2.2 Motion planning
We compare T3Former extensively with SOTA methods for autonomous driving, including LiDAR-based (IL [34], NMP [54], FF [13], EO [22]), camera-based (ST-P3 [14], UniAD [16], VAD [19], OccNet [40]), and occupancy-based methods (OccNet [40], OccWorld [58], RenderWorld [53], OccLLaMA [46]). The results are shown in Table 2.
@@ 4.2.2 Motion planning
Specifically, LiDAR- and camera-based methods explore various combinations of auxiliary supervisions to enhance the quality of planned motion trajectories. Despite their success, the auxiliary information they rely on is labor-intensive to obtain. In contrast, occupancy-based methods generate motion proposals using only occupancy representations. T3Former achieves the lowest L2 error over time, the lowest collision rate at the final time step, and the lowest average collision rate across all future motions. Compared to SOTA occupancy-based methods, T3Former shows excellent short-term and long-term planning (0s and 3s). This is primarily because we can establish precise, task-relevant correlations between scene changes and motion trajectories in a deep latent space, effectively filtering out noise from scene variations.
Figure 6: Visualization of 4D occupancy forecasting and motion planning for the next 3 seconds on the Occ3D test dataset (zoom in for a clearer view).